# Debt Behind the AI Boom: A Large-Scale Empirical Study of AI-Generated Code in the Wild

### 30 Mar 2026 v1

@@ 3.2. Commit-Level Quality Analysis
$3.$ Debt Lifecycle Analysis
@@
Issue Survival.  For each issue introduced by an AI-authored commit, we check whether it still exists at the repository’s latest revision (i.e., HEAD). If the file has been renamed, we follow its history using git log --follow. We then run static analysis on the corresponding file at HEAD. Next, we look for the same issue in the analysis results. We do not rely on the line number alone, since the location of the issue may move as the file changes. Instead, we match issues using their rule identifier together with a small amount of surrounding code context. If a match is found, the issue is classified as surviving. Otherwise, it is classified as not surviving.
@@ 4. Experimental Setup
Table 1. Summary of AI-authored commits by coding tool.
@@
GitHub Copilot 117,851 4,473 26.3
Claude 139,300 2,431 57.3
Cursor 19791 717 27.6
Gemini 12770 681 18.8
Devin 14650 300 48.8
Total 304,362 6,275* 48.5
*Repositories may use more than one tool; the total is deduplicated.
4.1. Dataset Summary
We collected our dataset by mining GitHub and applying filtering criteria. In our study, we focus on repositories containing production source files in Python, JavaScript, or TypeScript. Also, each repository must have at least one AI-authored commit. After filtering, we found 6,699 GitHub repositories with AI-authored commits, covering 29 AI coding tools. However, some tools have few commits, which may not provide reliable data for comparison. Thus, we focus on the five assistants with more than 10,000 attributed commits: GitHub Copilot, Claude, Cursor, Gemini, and Devin. Our final dataset includes 6,275 public GitHub repositories with 304,362 verified AI-authored commits. Table 1 summarizes the distribution of commits across the five AI coding assistants in our dataset. Figure 5(a) shows the monthly growth of AI-authored commits, with a sharp increase starting from mid-2025. Figure 5(b) shows that our dataset covers repositories with a wide range of popularity levels.
@@ 4.1. Dataset Summary
Figure 5. Overview of our dataset: (a) growth of AI-authored commits over time, and (b) distribution of repositories by GitHub star count as of March 2026.
4.2. Research Questions
@@
RQ1: What kinds of technical debt are introduced by AI coding assistants? This question investigates the basic characteristics of AI-introduced technical debt. We study what types of debt they introduce, including code smells and security issues. We also analyze how these issues are distributed across severity levels, languages, and rules.
@@
RQ3: How does AI-introduced technical debt evolve over time? Introducing debt is not necessarily a problem if it gets fixed quickly. The real concern is debt that persists unnoticed. We study whether introduced issues remain in the latest version of the repository or disappear over time. We also examine whether the associated code and files are later modified, fixed, or refactored.
4.3. Evaluation Metrics
@@
For the debt lifecycle (RQ3), we use three metrics. First, we compute the net impact by comparing the number of issues introduced and fixed by AI-authored commits. Second, we measure the survival rate of introduced issues:
@@
Third, to normalize for differences in commit volume across time periods, we report the number of surviving issues per 100 AI-authored commits.
5. Results
This section presents the empirical results of our study and answers the three research questions raised in Section 4.2.
Table 2. Overview of AI-introduced technical debt by issue type.
Type # Introduced # Repos # Commits % of Total
Code Smells 431,850 3,789 25,895 89.1%
Runtime Bugs 28,149 663 1,646 5.8%
Security Issues 24,607 1,038 4,158 5.1%
All Issues 484,606 3,841 26,564 -
5.1. RQ1: Types and Patterns of AI-Introduced Debt
Overview.  Table 2 presents a summary of the technical debt introduced by AI coding assistants in our dataset. In total, we identified 484,606 introduced issues across 3,841 repositories (61.2% of 6,275 repositories) and 26,564 commits (8.7% of 304,362 commits). This shows that a non-trivial portion of AI-authored commits introduce quality issues, and that these issues affect a large number of real-world repositories. Among all introduced issues, code smells, runtime bugs, and security issues are the three main categories. Code smells are by far the most common, accounting for 89.1% of all introduced issues. Below, we discuss each type in detail with real-world examples.
Table 3. Top 5 most frequent rules violated by AI coding assistants for each issue type.
@@ 5.1. RQ1: Types and Patterns of AI-Introduced Debt
Code Smells Broad exception handling 41,723 8.6%
Unused variables or parameters 28,718 5.9%
Unused argument 24,444 5.0%
Shadowed outer variable 20,251 4.2%
Access to protected member 19,835 4.1%
Runtime Bugs Undefined variable or reference 23,091 4.8%
Redeclared symbol 1,870 0.4%
Possibly used before assignment 1584 0.3%
@@
Unsubscriptable object 130 0.0%
Security Issues Subprocess Without Shell Check 4,334 0.9%
Try-Except-Pass 4,040 0.8%
Partial Executable Path 2,539 0.5%
Insecure Random Generator 2,355 0.5%
Subprocess Import 1,718 0.4%
ArchiveBox/ArchiveBox (archivebox, 2025c), core/models.py:L594:598, d360798
@@
Listing 1: Code smell: broad exception handling and missing file encoding in ArchiveBox (Claude Code).
@@
Code Smells.  Code smells are maintainability problems that make code harder to understand, debug, and evolve (Fowler, 2018). They increase long-term maintenance costs, even if they do not cause immediate failures. This finding is consistent with prior work under controlled settings (Liu et al., 2024; Siddiq et al., 2024), but our study confirms that the same pattern also appears in real-world repositories. Table 3 lists the top 5 most common code smell patterns (e.g., broad exception handling, unused variables or parameters). These issues are often small and easy to overlook during code review. Listing 1 shows an example from ArchiveBox (archivebox, 2025a) (>>27k stars). In commit d36079829bed (archivebox, 2025c), Claude Code updated the metadata loading logic in ArchiveBox. But the new code introduces two code smells. First, the bare except: pass block catches all exceptions silently. This makes errors harder to detect and debug. Second, the open() function does not specify a file encoding. This can lead to inconsistent behavior across different platforms and locales, since the default encoding may vary (Naoki, 2021). These issues may not cause immediate failures, but they can lead to maintenance challenges and subtle bugs in the future.
firecrawl/firecrawl (firecrawl, 2025b), python-sdk/firecrawl/firecrawl.py:L4004:4047, fb99747
@@
Listing 2: Runtime bug: undefined variable causing NameError in Firecrawl (Devin).
@@
Runtime Bugs.  Runtime bugs are code defects that can cause the program to fail during execution. Compared with code smells, they are less frequent. From Table 2, 28,149 runtime bugs are identified, which cover 663 repositories and 1,646 commits. However, their impact is more direct and severe than code smells. Table 3 shows the top 5 most common runtime bugs, which include undefined variable or reference, redeclared symbol, access to member before definition, possibly used before assignment, and unsubscriptable object. These patterns suggest that AI-generated code may look locally correct, but still fail to stay consistent with the surrounding context. What is interesting in this table is that we identified 23,091 cases of undefined variable or reference. Listing 2 presents one such case from firecrawl (>>98k stars). In commit fb99747ba978 (firecrawl, 2025b), Devin added a call that passes cache=cache as an argument. However, cache is never defined in the method, which leads to a NameError when that path is executed. The maintainer later fixed the bug by removing the undefined argument (firecrawl, 2025a). This example shows that AI-generated code can introduce real runtime errors. These errors require additional human effort to fix later.
microsoft/data-formulator (Microsoft, 2025), /tables_routes.py:L881:889, d8549c0
@@
Listing 3: Security issue: possible SQL injection in data-formulator (Copilot).
@@
Security Issues.  Security issues are another concern in AI-generated code. They pose a threat to data privacy and system safety. If not properly addressed, they can lead to data breaches or unauthorized access. Thus, it is important to identify and fix security issues early before they enter production repositories. As shown in Table 2, potentially insecure code patterns are also detected in 1,038 repositories and 4,158 commits. Table 3 shows that the most common security issues include subprocess without shell check, try-except-pass, insecure random generator, partial executable path, and unsafe HuggingFace Hub download. These patterns suggest that AI-generated code can introduce unsafe practices in process execution, error handling, and external resource usage. Figure 2 shows an example of the most common issues (i.e., subprocess without shell check) from hysteria2 (>>1.5k stars). A Copilot-authored commit (seagullz4, 2025b) introduced a shell=True subprocess call, which can enable command injection if user input reaches the call. A human developer later identified and removed the unsafe flag (seagullz4, 2025a). Beyond that, Listing 3 shows another example of a security issue (SQL injection) in data-formulator (>>1.2k stars). This repository is developed by Microsoft and uses GitHub Copilot for code generation. In commit d8549c0 (Microsoft, 2025), Copilot added a backend endpoint that constructs a SQL query by directly interpolating a user-supplied table name. This creates a potential SQL injection vector. If an attacker can control the source_name variable, they can inject malicious SQL code that may lead to data breaches or unauthorized access. This issue remained in the repository for several weeks before the maintainer refactored the code and removed the unsafe SQL construction (Microsoft, 2026).
@@
Table 4. Top 5 most frequent rules by language. Rate indicates each rule’s share of all issues in that language.
@@
Python Broad exception handling 41,721 14.1%
Unused argument 24,444 8.3%
Undefined variable or reference 23,091 7.8%
Access to protected member 19,835 6.7%
Unused import 17,435 5.9%
JavaScript/ TypeScript Unused variables or parameters 28,718 15.2%
Shadowed outer variable 18,016 9.6%
Block-scoped variable misuse 11255 6.0%
No sequences $89$ 4.8%
No unused expressions 7673 4.1%
Programming Language Differences.  Table 4 compares the top 5 most frequent rules in Python and JavaScript/TypeScript. There are some patterns that are common in both languages. For example, both languages show issues related to unused code (i.e., unused arguments in Python, unused variables in JavaScript/TypeScript). At the same time, each language also has its own characteristic issues. Python’s top rules are dominated by exception handling and dynamic typing problems, while JavaScript/TypeScript issues tend to involve scoping and variable declaration patterns. This observation is consistent with prior studies (Liu et al., 2024; Siddiq et al., 2024), which also found that the types of issues in AI-generated code can vary depending on the programming language and the tools used. These results suggest that some debt patterns may be language-specific. However, the overall trend (e.g., code smells dominate) holds across both languages.
Answer to RQ1: AI-generated code introduces technical debt in the form of code smells (89.1%), runtime bugs (5.8%), and security issues (5.1%). Among them, code smells are by far the most common.
Figure 6. Percentage of commits with issues and total commit volume per AI coding assistant.
Table 5. Average number of issues introduced per commit by type. Darker cells indicate higher values.
5.2. RQ2: Comparison Across AI Coding Assistants
In this RQ, we examine how technical debt patterns vary across AI coding assistants. Figure 6 shows the percentage of commits with issues for each of the five AI coding assistants. What stands out in this figure is that more than 15% of commits by each AI tool introduce at least one issue. The rates also vary across tools, ranging from 17.3% for GitHub Copilot to 28.7% for Gemini. However, technical debt does not disappear when using different or newer tools. This suggests that technical debt is a common and persistent problem in AI-generated code.
Table 5 further compares the average number of introduced issues per commit by type. We can see that all five tools share a common pattern, where the code smell rate is much higher than the bug and security issue rates. This is consistent with our findings in RQ1. At the same time, there are also differences across tools. From Table 5, we can see that Claude has the highest issue rate per commit (1.96), while Devin has the lowest (0.87). These differences may be due to differences in usage patterns and development context, rather than the tools alone. Still, it is apparent that the overall pattern of technical debt is consistent across all five tools.

### 26 Apr 2026 v2

@@ III-B Commit-Level Quality Analysis
III-C Debt Lifecycle Analysis
@@
Issue Survival.  For each issue introduced by an AI-authored commit, we check whether it still exists at the repository’s latest revision (i.e., HEAD). If the file has been renamed, we follow its history using git log --follow. We then run static analysis on the corresponding file at HEAD. Next, we look for the same issue in the analysis results. We do not rely on the line number alone, since the location of the issue may move as the file changes. Instead, we match issues using their rule identifier together with a small amount of surrounding code context. If a match is found, the issue is classified as surviving. Otherwise, it is classified as not surviving. In other words, an introduced issue is counted as surviving only if the same issue is still present at HEAD. If the original issue disappears and a different issue appears later, the original issue is treated as not surviving.
@@ IV Experimental Setup
TABLE I: Summary of AI-authored commits by coding tool.
@@
GitHub Copilot 118,012 4,487 26.3
Claude 138,249 2,434 56.8
Cursor 19,587 714 27.4
Gemini 12,429 681 18.3
Devin 14,302 298 48.0
Total 302,579 6,299* 48.0
Each name groups all variants of the tool (e.g., Claude includes Claude Code and its different model versions). *Repositories may use more than one tool; the total is deduplicated.
IV-A Dataset Summary
We collected our dataset by mining GitHub and applying filtering criteria described in Section III-A. After repository-level filtering, we obtained 6,699 repositories with confirmed AI-authored commits, covering 29 AI coding tools. However, some tools have very few commits, which may not provide reliable data for comparison. Thus, we focus on the five assistants with more than 10,000 attributed commits: GitHub Copilot [24], Claude [1], Cursor [10], Gemini [26], and Devin [7]. This results in 6,412 repositories with 317.4K AI-attributed commits. For the commit-level quality analysis, we further exclude cases that cannot be analyzed reliably such as repositories that became unavailable, deletion-only commits, and commits that do not modify production Python, JavaScript, or TypeScript source files. Our final analysis dataset therefore includes 6,299 public GitHub repositories with 302.6K analyzed AI-attributed commits. Table I summarizes the distribution of commits across the five AI coding assistants in our dataset. Figure 6(a) shows the monthly growth of AI-authored commits, with a sharp increase starting from mid-2025. Figure 6(b) shows that our dataset covers repositories with a wide range of popularity levels.
@@ IV-A Dataset Summary
Figure 6: Overview of our dataset: (a) growth of AI-authored commits over time, and (b) distribution of repositories by GitHub star count as of March 2026.
IV-B Research Questions
@@
RQ1: What kinds of technical debt are introduced by AI coding assistants? This question investigates the basic characteristics of AI-introduced technical debt. We study what types of debt they introduce, including code smells, correctness issues, and security issues. We also analyze how these issues are distributed across languages, rules, and AI coding assistants.
@@
RQ3: To what extent does AI-introduced technical debt persist in the codebase? Introducing debt is not necessarily a problem if it gets fixed quickly. The real concern is debt that persists unnoticed. We compare the number of issues introduced and fixed by AI-authored commits to assess the net impact. We study whether introduced issues remain in the latest version of the repository or disappear over time.
IV-C Evaluation Metrics
@@
For the debt lifecycle (RQ3), we use two metrics. First, we compute the net impact by comparing the number of issues introduced and fixed by AI-authored commits. Second, we measure the survival rate of introduced issues:
@@
IV-D Validation
To assess the reliability of our pipeline, two authors independently inspected random samples of AI-attributed commits and detected issues.
We randomly sampled 100 AI-attributed commits. For each commit, we manually verified whether it was correctly attributed to the claimed AI coding assistant using the Git metadata and commit message described in Section III-A. One sampled case could not be verified because the corresponding repository was no longer available. For the remaining 99 verifiable commits, both authors independently labeled all cases, and all 99 were confirmed as correctly attributed after manual inspection. This yields a conservative attribution precision of 99.0% over the full sample, or 100% over the verifiable subset.
We also randomly sampled 100 introduced issues from manually verified AI-attributed commits and validated two aspects of each: (1) whether the reported issue was a real issue, rather than a false positive from the static analysis tools; and (2) whether the survival classification at HEAD was correct (i.e., whether the issue was correctly labeled as surviving or not surviving). After excluding one case with unavailable or incomplete validation context, both authors independently labeled 99 issue cases. For issue validity, the raw agreement was 95/99 = 0.960, with Cohen’s $\kappa=0.851$, indicating almost perfect agreement. For survival classification, the raw agreement was 97/99 = 0.980, with Cohen’s $\kappa=0.960$, also indicating almost perfect agreement.
Using the adjudicated labels as reference, the pipeline achieved an accuracy of 85.9%, precision of 85.9%, recall of 100.0%, and F1 of 92.4% for issue validity. For survival classification, the pipeline achieved an accuracy of 84.8%, precision of 86.7%, recall of 81.2%, and F1 of 83.9%. These results suggest that our pipeline provides reliable large-scale estimates, while still leaving some room for noise in issue detection and lifecycle tracking.
V Results
This section presents the empirical results of our study and answers the three research questions raised in Section IV-B.
TABLE II: Overview of AI-introduced technical debt by issue type.
Type # Introduced # Repos # Commits % of Issues
Code Smells 432,748 3,812 25,740 89.3%
Correctness Issues 28,931 665 1,650 6.0%
Security Issues 22,687 1,643 5,142 4.7%
All Issues 484,366 3,946 27,677 -
V-A RQ1: Types and Patterns of AI-Introduced Debt
Overview.  Table II presents a summary of the technical debt introduced by AI coding assistants in our dataset. In total, we identified 484,366 introduced issues across 3,946 repositories (62.6% of 6,299 repositories) and 27,677 commits (9.1% of 302,579 commits). This shows that a non-trivial portion of AI-authored commits introduce quality issues, and that these issues affect a large number of real-world repositories. Among all introduced issues, code smells, correctness issues, and security issues are the three main categories. Code smells are by far the most common, accounting for 89.3% of all introduced issues. Below, we discuss each type in detail with real-world examples.
TABLE III: Top 5 most frequent rules violated by AI coding assistants for each issue type.
@@ V-A RQ1: Types and Patterns of AI-Introduced Debt
Code Smells Broad exception handling 41,374 8.5%
Unused variables or parameters 28,272 5.8%
Unused argument 24,357 5.0%
Shadowed outer variable 20,647 4.3%
Access to protected member 19,796 4.1%
Correctness Issues Undefined variable or reference 23,856 4.9%
Redeclared symbol 1,888 0.4%
Possibly used before assignment 1575 0.3%
@@
Unsubscriptable object 176 0.0%
Security Issues Path traversal via path.join/resolve 8,677 1.8%
Unsafe format string 4,792 1.0%
Non-literal regular expression 1,212 0.3%
Child process execution 607 0.1%
SQLAlchemy raw query execution 591 0.1%
ArchiveBox [5], core/models.py:L594:598, d360798
@@
Listing 1: Code smell: broad exception handling and missing file encoding in ArchiveBox [3] (Claude Code).
@@
Code Smells.  Code smells are maintainability problems that make code harder to understand, debug, and evolve [19]. They increase long-term maintenance costs, even if they do not cause immediate failures. This finding is consistent with prior work under controlled settings [35, 62], but our study confirms that the same pattern also appears in real-world repositories. Table III lists the top 5 most common code smell patterns (e.g., broad exception handling, unused variables or parameters). These issues are often small and easy to overlook during code review. Listing 1 shows an example from ArchiveBox [3] (>>27k stars). In commit d36079829bed [5], Claude Code updated the metadata loading logic in ArchiveBox. But the new code introduces two code smells. First, the bare except: pass block catches all exceptions silently. This makes errors harder to detect and debug. Second, the open() function does not specify a file encoding. This can lead to inconsistent behavior across different platforms and locales, since the default encoding may vary [43]. These issues may not cause immediate failures, but they can lead to maintenance challenges and subtle bugs in the future.
firecrawl [17], firecrawl.py:L4004:4047, fb99747
@@
Listing 2: Correctness issues: undefined variable causing NameError in Firecrawl [18] (Devin).
@@
Correctness Issues.  Correctness issues are code defects that can cause the program to fail during execution. Compared with code smells, they are less frequent. From Table II, 28,931 correctness issues are identified, which cover 665 repositories and 1,650 commits. However, their impact is more direct and severe than code smells. Table III shows the top 5 most common correctness issues, which include undefined variable or reference, redeclared symbol, access to member before definition, possibly used before assignment, and unsubscriptable object. These patterns suggest that AI-generated code may look locally correct, but still fail to stay consistent with the surrounding context. What is interesting in this table is that we identified 23,856 cases of undefined variable or reference. Listing 2 presents one such case from firecrawl (>>98k stars). In commit fb99747ba978 [17], Devin added a call that passes cache=cache as an argument. However, cache is never defined in the method, which leads to a NameError when that path is executed. The maintainer later fixed the bug by removing the undefined argument [16]. This example shows that AI-generated code can introduce real correctness issues. These errors require additional human effort to fix later.
data-formulator [39], tables_routes.py:L881:889, d8549c0
@@
Listing 3: Security issue: possible SQL injection in microsoft/data-formulator [40] (Copilot).
@@
Security Issues.  Security issues are another concern in AI-generated code. In our study, this category includes not only direct security vulnerabilities, but also insecure coding patterns that can be viewed as security debt. Some of these issues may be exploitable at the time they are introduced, while others may become security risks after later code changes or broader system integration. Thus, it is important to identify and fix these issues early before they accumulate in production repositories. As shown in Table II, potentially insecure code patterns are detected in 1,643 repositories and 5,142 commits. Table III shows that common security issues such as path traversal via path.join or path.resolve, unsafe format strings, non-literal regular expressions, and child process execution. These patterns suggest that AI-generated code can introduce unsafe practices in process execution, file path handling, and string formatting. A common pattern across these issues is unsafe handling of untrusted input, where user- or context-controlled values flow into security-sensitive operations without proper validation or sanitization. Figure 2 shows one such example from hysteria2 (>>1.5k stars). A Copilot-authored commit [59] introduced a shell=True subprocess call. This pattern is not necessarily an exploitable vulnerability by itself in the current version, but it creates security debt because it can enable command injection if untrusted input later reaches the command string. A human developer later identified and removed the unsafe flag [58]. Beyond that, Listing 3 shows another example of a security issue (SQL injection) in data-formulator (>>1.2k stars). This repository is developed by Microsoft and uses GitHub Copilot for code generation. In commit d8549c0 [39], Copilot added a backend endpoint that constructs a SQL query by directly interpolating a user-supplied table name. This creates a potential SQL injection vector. If an attacker can control the source_name variable, they can inject malicious SQL code that may lead to data breaches or unauthorized access. This issue remained in the repository for several weeks before the maintainer refactored the code and removed the unsafe SQL construction [40].
@@
TABLE IV: Top 5 most frequent rules by language. Rate indicates each rule’s share of all issues in that language.
@@
Python Broad exception handling 41,374 14.9%
Unused argument 24,357 8.8%
Undefined variable or reference 23,856 8.6%
Access to protected member 19,796 7.1%
Unused import 17,376 6.3%
JavaScript/ TypeScript Unused variables or parameters 28,272 13.6%
Shadowed outer variable 18,417 8.9%
Block-scoped variable misuse 11568 5.6%
No sequences 9358 4.5%
Path traversal via path.join/resolve 8677 4.2%
Programming Language Differences.  Table IV compares the top 5 most frequent rules in Python and JavaScript/TypeScript. There are some patterns that are common in both languages. For example, both languages show issues related to unused code (i.e., unused arguments in Python, unused variables in JavaScript/TypeScript). At the same time, each language also has its own characteristic issues. Python’s top rules are dominated by exception handling and dynamic typing problems, while JavaScript/TypeScript issues tend to involve scoping and variable declaration patterns. This observation is consistent with prior studies [35, 62], which also found that the types of issues in AI-generated code can vary depending on the programming language and the tools used. These results suggest that some debt patterns may be language-specific. However, the overall trend (e.g., code smells dominate) holds across both languages.
Answer to RQ1: AI-generated code introduces technical debt in the form of code smells (89.3%), correctness issues (6.0%), and security issues (4.7%). Among them, code smells are by far the most common.
Figure 7: Percentage of commits with issues and total commit volume per AI coding assistant.
TABLE V: Average number of issues introduced per commit by type. Darker cells indicate higher values.
V-B RQ2: Comparison Across AI Coding Assistants
In this RQ, we examine how technical debt patterns vary across AI coding assistants. Figure 7 shows the percentage of commits with issues for each of the five AI coding assistants. What stands out in this figure is that more than 15% of commits by each AI coding tool introduce at least one issue. The rates also vary across tools, ranging from 17.4% for GitHub Copilot to 29.1% for Gemini. This suggests that technical debt appears across all studied tools, although the rate differs by tool. In other words, the problem is not limited to a single AI coding assistant.
Table V further compares the average number of introduced issues per commit by type. We can see that all five tools share a common pattern, where the code smell rate is much higher than the correctness and security issue rates. This is consistent with our findings in RQ1. At the same time, there are also differences across tools. From Table V, we can see that Claude has the highest issue rate per commit (1.95), while Devin has the lowest (0.89). These differences may be due to differences in usage patterns and development context, rather than the tools alone. Still, it is apparent that the overall pattern of technical debt is consistent across all five tools.